# Numerical methods for solving PIDEs arising in swing option pricing under a two-factor mean-reverting model with jumps

### v1

@@ Abstract
This paper concerns the numerical valuation of swing options with discrete action times under a linear two-factor mean-reverting model with jumps. The resulting sequence of two-dimensional partial integro-differential equations (PIDEs) are convection-dominated and possess a nonlocal integral term due to the presence of jumps. Further, the initial function is nonsmooth. We propose various second-order numerical methods that can adequately handle these challenging features. The stability and convergence of these numerical methods are analysed theoretically. By ample numerical experiments, we confirm their second-order convergence behaviour.
@@ 1 Introduction
The electricity price possesses a mean reversion property with spikes and (daily, weekly and annually) seasonal patterns. One of the first models for this price was proposed by Lucia and Schwartz (2002) in the form of a geometric Ornstein–Uhlenbeck process which has a mean-reverting property. However, this model has the disadvantage of not incorporating spikes. In Benth et al. (2007), the electricity spot price is modelled as a linear combination of an Ornstein–Uhlenbeck process and pure mean-reverting jump processes. Hambly et al. (2009) considered an exponential form of this model. In this paper, we follow the approach of Eriksson et al. (2014) and consider an affine two-factor model with finite activity jumps. The choice of an affine model is motivated by the fact that the electricity market exhibits negative prices, especially in the recent years.
@@
This paper is organised in the following way. Section 2 presents the electricity spot price model under consideration and the formulation of the option pricing problem as a sequence of two-dimensional PIDEs. Section 3 concerns the spatial discretisation with special attention for the convection and integral parts. Section 4 presents the temporal discretisation schemes. These schemes are all second-order and treat the integral term in an explicit way by means of a fixed-point iteration. Section 5 is devoted to the theoretical analysis of the stability and convergence properties of the schemes. In Section 6, we present ample numerical experiments, especially to study the observed order of convergence. Finally, Section 7 gives our conclusions.
@@ 2.1 The electricity spot price
$S_{t}$ $=X_{t}+Y_{t},$ (2.1)
$dX_{t}$ $=\alpha(\mu-X_{t})dt+\sigma dW_{t},$ (2.2)
$dY_{t}$ $=-\beta Y_{t}dt+J_{t}dN_{\lambda,t}.$ (2.3)
@@
The electricity spot price could be modelled in exponential form, i.e., $S_{t}=\exp(X_{t}+Y_{t})$, see, e.g., Hambly et al. (2009), Calvo-Garrido et al. (2016, 2019). However, we prefer the affine form, i.e., $S_{t}=X_{t}+Y_{t}$, which allows the price to become negative. The occurrence of negative prices was observed in the day-ahead market and it is generally due to oversupply combined with low demand, inflexible power stations (e.g., nuclear reactors) and cheap renewable power. A study by the Belgian Federal Commission for Electricity and Gas Regulation (CREG), see Tirez et al. (2023), indicated that, in 2020, the cumulative number of hours with negative prices reached 136 hours in Belgium, 102 hours in France, 319 hours in Germany and 97 hours in the Netherlands.
@@ 2.2 Formulation of the problem
    Swing action times are in the form $T_{n}=n\Delta T$ ($n=1,\ldots,N_{a}$) with $\Delta T=\frac{T}{N_{a}}$.
@@ 3.1 Diffusion-reaction part
The spatial domain $\mathbb{R}^{2}$ is truncated to a bounded set $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$, where $x_{\rm max}>0$ and $y_{\rm max}>0$ and $x_{\rm min}<0$ and $y_{\rm min}<0$ are all taken sufficiently large in absolute value. For the $x$-direction and $y$-direction, we impose linear boundary conditions:
@@ 3.3 Convection part
Aside from the nonlocal integral term in (2.5), a significant challenge arises due to the nature of the asset price model: the problem is convection-dominated in the $x$-direction and exhibits pure convection in the $y$-direction. This feature is attributed to the electricity price dynamics, which exhibits fast mean reversion characterised by large values of the parameters $\alpha$ and, especially, $\beta$. These values, detailed in the parameter sets provided in Table 1 of Section 6.1, result in a highly convection-dominated problem as the diffusion coefficient, $\frac{\sigma^{2}}{2}$, is constant and relatively small in comparison to the convection coefficients.
@@ 4.2 Temporal schemes for the semidiscretisation approach
This scheme, studied by in ’t Hout (2025), combines the DIRK scheme with a penalty/fixed-point iteration for the numerical valuation for American-style options under the two-asset Kou-type jump-diffusion model. In our case, we obtain.
@@ 5.2 Stability and convergence study
$\big(I-\tfrac{1}{2}\Delta tA\big)Y_{\ell}=(I+\tfrac{1}{2}\Delta tA)V^{n}+\tfrac{1}{2}\Delta t(BV^{n}+BY_{\ell-1})+\tfrac{1}{2}\Delta t(g_{n}+g_{n+1}).$ (5.10)
@@
denotes the local spatial error.
By the smoothness assumption on $v_{\Delta x,\Delta y}$, Taylor expansion yields for some positive constant $C_{1}$ (independent of $i$, $\Delta t$, $\Delta x$, $\Delta y$) that
$\|v_{\Delta x,\Delta y}(t_{i+1})-v_{\Delta x,\Delta y}(t_{i})-\tfrac{1}{2}\Delta t\Bigl(\frac{dv_{\Delta x,\Delta y}}{dt}(t_{i+1})+\frac{dv_{\Delta x,\Delta y}}{dt}(t_{i})\Bigr)\|_{2}\leq C_{1}\Delta t^{3}.$
The semidiscretisation is consistent of second-order, i.e., by definition there exists a positive constant $C_{2}$ (independent of $i$, $\Delta x$, $\Delta y$) such that
$\delta(t_{i+1})||_{2}<C_{2}(\Delta x^{2}+\Delta y^{2}).$
@@
From Lemmas 5.2 and 5.4, we have: $\mu_{2}[\tfrac{1}{2}\Delta t(A+B)]\leq\tfrac{1}{2}\Delta t\widehat{C}$. Thus, if $\widehat{C}\Delta t\leq 1$, then
@@ 6 Numerical experiments
In this section, we present a series of numerical experiments for the case of European call and swing options. The main objective is to experimentally validate the proposed numerical schemes. For the European call option, we investigate the convergence behaviour of the total and temporal errors, see Section 6.2, which gives us an indication of the convergence behaviour for swing options. The study for swing options will be focused on the convergence behaviour of the temporal error, see Section 6.4. The parameter sets employed in the numerical experiments are detailed in Section 6.1.
We apply cell averaging in the definition of the initial vector $V(0)=V^{0}$ corresponding to the option payoff, because relying fully on its pointwise evaluation can lead to a deteriorated spatial convergence behaviour. Let
@@ 6.1 Financial parameter values
Table 2: Parameter sets with Kou-type jump case. The time is measured in years.
@@ 6.2 Convergence behaviour: European call option
Figures 2 and 3 show that, for all schemes and all parameters sets, the total error decreases monotonically as $N$ and $m$ increase in a directly proportional way. For the parameter sets 1, 2, 4, 5, a favourable second-order convergence behaviour is observed. Further, for each of these four sets, CNFI and DIRKFI are seen to have about the same error constants, which is always smaller than that for SLCNFI, and hence, CNFI and DIRKFI are to be preferred over SLCNFI.
The sets 3 and 6 represent highly convection-dominated problems as the volatility $\sigma$ is small and the jump intensity $\lambda$ is low. These characteristics lead to the emergence of a region of nonsmoothness where the solution $v$ has steep gradients. In this situation, the convergence order for the total error of CNFI and DIRKFI reduces: it is (asymptotically) equal to 1.6. On the other hand, for SLCNFI the convergence order remains (asymptotically) equal to two. Further, its error constant is smaller than that for CNFI and DIRKFI. Hence, for these two sets, SLCNFI is to preferred. We remark, however, that sets 3 and 6 are less representative of realistic market situations, since electricity prices typically experience significant fluctuations and frequent jumps, which are not captured well by a small volatility and low jump intensity.
@@
Figure 3: European call option in the Kou-type jump model. Total discretisation errors of the SLCNFI, CNFI and DIRKFI schemes for $N=\left\lceil{\frac{m}{2}}\right\rceil$ and set 1 (top), set 2 (middle), set 3 (bottom). Added: dashed reference line for convergence order 2.
@@
Figure 5: European call option in the Kou-type jump model. Temporal discretisation errors of the CNFI and DIRKFI schemes for $m=200$ and set 1 (top), set 2 (middle), set 3 (bottom). Added: dashed reference line for convergence order 2.
@@ 6.3 Numerical valuation of swing options
For the numerical valuation of swing options, we combine the numerical method proposed for the European call option case with the dynamic programming principle to find the optimal exercise at each action time. As before, we reverse the time to obtain an initial condition instead of a terminal condition. The action time $\tau_{n}$ in reversed time is $T-T_{N_{a}-n+1}$ ($n=1,2,\ldots,N_{a}$) and the sequence of PIDEs (2.4) then becomes
@@
The valuation procedure is outlined in Algorithm 1. At the first action time, $\tau_{1}=0$, the holder buys the maximal allowable amount. Given the option value at $\tau_{1}$, the PIDE is solved to obtain the option value up to the second action time, $\tau_{2}$. It is then considered whether it is optimal to exercise or not by maximizing the option value at $\tau_{2}$. This process is repeated for each subsequent action time up to and including $\tau_{N_{a}}$. At each action time, the option value is determined by optimizing over all feasible exercise amounts $b\in\{0,1,\ldots,L\}$, while ensuring that the cumulative purchased amount does not exceed the global constraint $M$. Finally, the PIDE is solved over the interval $(\tau_{N_{a}},\tau_{N_{a}+1}]$ to arrive at the desired option value function $v(\cdot,\cdot,0,T)$.
@@
Figure 6 displays the graph of the swing option value function for $z=0$ at $t=T$ (reversed time). Here the number of swing action times is set to $N_{a}=20$, with local constraint specified by $L=1$ and global constraint by $M=10$. Both Merton- and Kou-type jump models are considered, for the parameter sets 1 and 4 given by the Tables 1 and 2, respectively, with maturity time $T=1$. Table 3 provides corresponding reference values obtained with the CNFI scheme and $N=100$ time steps between each two successive action times.
Figure 6: Swing option value functions with $z=0$ and $t=T=1$ for set 1 (top) and set 4 (bottom) where $N_{a}=20$, $L=1$, $M=10$.
Table 3: Swing option values with $z=0$ and $t=T=1$ for sets 1 and 4 where $N_{a}=20$, $L=1$, $M=10$ and using the CNFI scheme $m_{1}=m_{2}=N=100$.
@@
$x=40$ $x=60$ $x=80$
$y=5$ $686.5660$ $699.2691$ $714.7266$
y=100y=-100 $686.5246$ $699.2137$ $714.6424$
$y=100$ $686.6036$ $699.3196$ $714.8036$
6.4 Convergence behaviour: swing options
In this section, we numerically study the temporal discretisation error for the CNFI and DIRKFI schemes in the context of swing options. Here again $N_{a}=20$, $L=1$, $M=10$ and the parameter sets 1 and 4 are considered with $T=1$.
Figure 7: Swing option with $N_{a}=20,\,L=1,\,M=10$. Temporal discretisation errors of the CNFI and DIRKFI schemes for $m=100$ and set 1 (top) and set 4 (bottom). Added: dashed reference line for convergence order 2.
Figure 7 displays the temporal errors for 20 different values of $N$ ranging from 50 to 500 and $m=100$. Notice that the total number of time steps over $[0,T]$ then ranges from $10^{3}$ and $10^{4}$. A reference solution has been computed by applying the CNFI scheme with $N=4000$.
As in the European case (see Section 6.2), the temporal errors for both the CNFI and DIRKFI schemes decrease monotonically as the time step is refined. Figure 7 reveals that both methods exhibit a favourable, second-order convergence behaviour. Additional experiments with larger numbers of spatial grid points ($m=200,300$) further show that the temporal errors are essentially unaffected, as desired.
We note that, unlike the European case, the temporal errors for CNFI and DIRKFI are nearly identical. This is attributed to the structure of the swing option, which involves discrete action times where the option price is updated, in addition to the continuation phase. These updates apparently reduce the distinction between the two schemes in terms of temporal accuracy.

### v2

@@ Abstract
This paper concerns the numerical valuation of swing options with discrete action times under a linear two-factor mean-reverting model with jumps. The resulting sequence comprises two-dimensional partial integro-differential equations (PIDEs); each equation is convection-dominated and possesses a nonlocal integral term due to the presence of jumps. Further, the initial function is nonsmooth. We propose various second-order numerical methods that can adequately handle these challenging features. The stability and convergence of these numerical methods are analysed theoretically. By ample numerical experiments, we confirm their second-order convergence behaviour. In addition to the numerical valuation of swing options, a study of their Delta Greeks and the optimal exercise policy is performed.
@@ 1 Introduction
The electricity price possesses a mean reversion property with spikes and (daily, weekly and annually) seasonal patterns. One of the first models for this price was proposed by Lucia and Schwartz (2002) in the form of a geometric Ornstein–Uhlenbeck process which has a mean-reverting property. However, this model has the disadvantage of not incorporating spikes. In Benth et al. (2007), the electricity spot price is modelled as a linear combination of an Ornstein–Uhlenbeck process and pure mean-reverting jump processes. Hambly et al. (2009) considered an exponential form of this model. In this paper, we follow the approach of Eriksson et al. (2014) and consider an affine two-factor model with finite activity jumps. The choice of an affine model is motivated by the fact that the electricity market exhibits negative prices, especially in recent years.
@@
The fourth contribution of this paper is the study of swing options using a linear two-factor model for the underlying electricity price. Using the proposed numerical methods, we focus first on a rigorous valuation of this type of options. Then, we study the impact of the model parameters on the option value and also the behaviour of the optimal exercise policy.
This paper is organised in the following way. Section 2 presents the electricity spot price model under consideration and the formulation of the option pricing problem as a sequence of two-dimensional PIDEs. Section 3 concerns the spatial discretisation with special attention to the convection and integral parts. Section 4 presents the temporal discretisation schemes. These schemes are all second-order and treat the integral term in an explicit way by means of a fixed-point iteration. Section 5 is devoted to the theoretical analysis of the stability and convergence properties of the schemes. In Section 6, we present ample numerical experiments, especially to study the observed order of convergence. In addition, the behaviour of the Delta Greeks and the optimal exercise policy are investigated. Finally, Section 7 gives our conclusions.
@@ 2.1 The electricity spot price
$S_{t}$ $=X_{t}+Y_{t},$
$dX_{t}$ $=\alpha(\mu-X_{t})dt+\sigma dW_{t},$
$dY_{t}$ $=-\beta Y_{t}dt+J_{t}dN_{\lambda,t}.$
@@
The electricity spot price could be modelled in exponential form, i.e., $S_{t}=\exp(X_{t}+Y_{t})$, see, e.g., Hambly et al. (2009), Calvo-Garrido et al. (2016, 2019). However, we prefer the affine form, i.e., $S_{t}=X_{t}+Y_{t}$, which allows the price to become negative. The occurrence of negative prices was observed in the day-ahead market and is generally due to oversupply combined with low demand, inflexible power stations (e.g. nuclear reactors) and cheap renewable power. A study by the Belgian Federal Commission for Electricity and Gas Regulation (CREG), see Tirez et al. (2023), indicated that, in 2020, the cumulative number of hours with negative prices reached 136 hours in Belgium, 102 hours in France, 319 hours in Germany and 97 hours in the Netherlands.
@@ 2.2 Formulation of the problem
    Swing action times are of the form $T_{n}=n\Delta T$ ($n=1,\ldots,N_{a}$) with $\Delta T=\frac{T}{N_{a}}$.
@@ 3.1 Diffusion-reaction part
The spatial domain $\mathbb{R}^{2}$ is truncated to a bounded set $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$, where $x_{\rm max}>0$, $y_{\rm max}>0$, $x_{\rm min}<0$ and $y_{\rm min}<0$ are all taken sufficiently large in absolute value. For the $x$-direction and $y$-direction, we impose linear boundary conditions:
@@ 3.3 Convection part
Aside from the nonlocal integral term in (2.2), a significant challenge arises due to the nature of the asset price model: the problem is convection-dominated in the $x$-direction and exhibits pure convection in the $y$-direction. This feature is attributed to the electricity price dynamics, which exhibits fast mean reversion characterised by large values of the parameters $\alpha$ and, especially, $\beta$. These values, detailed in the parameter sets provided in Table 1 of Section 6.1, result in a highly convection-dominated problem as the diffusion coefficient, $\frac{\sigma^{2}}{2}$, is constant and relatively small in comparison with the convection coefficients.
@@ 4.2 Temporal schemes for the semidiscretisation approach
This scheme, studied by in ’t Hout (2025), combines the DIRK scheme with a penalty/fixed-point iteration for the numerical valuation for American-style options under the two-asset Kou-type jump-diffusion model. In our case, we obtain
@@ 5.2 Stability and convergence study
$\big(I-\tfrac{1}{2}\Delta tA\big)Y_{\ell}=(I+\tfrac{1}{2}\Delta tA)V^{n}+\tfrac{1}{2}\Delta t(BV^{n}+BY_{\ell-1})+\tfrac{1}{2}\Delta t(g_{n}+g_{n+1}),$ (5.10)
@@
denotes the local spatial error. By assumption, the semidiscretisation is consistent of order two, i.e., there exists a positive constant $C_{1}$ (independent of $t$, $\Delta x$, $\Delta y$) such that
$\delta(t)||_{2}\leq C_{2}(\Delta x^{2}+\Delta y^{2}).$
Next, by the smoothness assumption on $v_{\Delta x,\Delta y}$, Taylor expansion yields for some positive constant $C_{2}$ (independent of $i$, $\Delta t$, $\Delta x$, $\Delta y$) that
$\|v_{\Delta x,\Delta y}(t_{i+1})-v_{\Delta x,\Delta y}(t_{i})-\tfrac{1}{2}\Delta t\Bigl(\frac{dv_{\Delta x,\Delta y}}{dt}(t_{i+1})+\frac{dv_{\Delta x,\Delta y}}{dt}(t_{i})\Bigr)\|_{2}\leq C_{2}\Delta t^{3}.$
@@
From Lemmas 5.2 and 5.4, we have $\mu_{2}[\tfrac{1}{2}\Delta t(A+B)]\leq\tfrac{1}{2}\Delta t\widehat{C}$. Thus, if $\widehat{C}\Delta t\leq 1$, then
@@
For the DIRK scheme (5.16) with $\theta\in[\frac{1}{4},\frac{1}{2}]$ the same second-order convergence result holds as Theorem 5.7 for the Crank–Nicolson scheme. A direct adaptation of the proof of Theorem 5.7 to the case of the DIRK scheme yields only first-order convergence, however (when $\theta\not=\frac{1}{2}$). Still, second-order convergence can be proved by invoking a key condition of (Hundsdorfer 1992, Lemma 5.1), see also (Hundsdorfer and Verwer 2003, Sect. II.2.3), which employs a special structure of the local error. For the sake of brevity, we omit the details of the proof of second-order convergence for the DIRK scheme in this paper. We mention that similar convergence proofs, using Hundsdorfer’s Lemma, have been given in, e.g., Hundsdorfer (2002) and in ’t Hout and Wyns (2016) for closely related (splitting) schemes.
@@ 6 Numerical experiments
In this section, we present a series of numerical experiments for European call options and swing options. A main objective is to experimentally validate the proposed numerical schemes. For the European call option, we investigate the convergence behaviour of the total and temporal errors, see Section 6.2, which gives us an indication of the convergence behaviour for swing options. The numerical valuation algorithm for swing options is described in Section 6.3. The subsequent study for swing options deals with the convergence behaviour of the total error, the Delta Greeks and the optimal exercise policy, see Sections 6.4, 6.5 and 6.6 respectively.
In the following, we apply cell averaging in the definition of the initial vector $V(0)=V^{0}$ corresponding to the option payoff, because relying fully on its pointwise evaluation can lead to a deteriorated spatial convergence behaviour. Let
@@ 6.1 Financial parameter values
Table 2: Parameter sets for Kou-type jump case. The time is measured in years.
@@ 6.2 Convergence behaviour: European call option
Figures 2 and 3 show that, for all schemes and all parameters sets, the total error decreases monotonically as $N$ and $m$ increase in a directly proportional way. For the parameter sets 1, 2, 4, 5, a favourable second-order convergence behaviour is observed. Further, for each of these four sets, CNFI and DIRKFI are seen to have about the same error constants, which are always smaller than that for SLCNFI, and hence, CNFI and DIRKFI are to be preferred over SLCNFI.
The sets 3 and 6 represent highly convection-dominated problems as the volatility $\sigma$ is small and the jump intensity $\lambda$ is low. These characteristics lead to the emergence of a region of nonsmoothness where the solution $v$ has steep gradients. In this situation, the convergence order for the total error of CNFI and DIRKFI reduces: it is (asymptotically) equal to 1.6. On the other hand, for SLCNFI the convergence order remains (asymptotically) equal to two. Further, its error constant is smaller than that for CNFI and DIRKFI. Hence, for these two sets, SLCNFI is preferred. We remark, however, that sets 3 and 6 are less representative of realistic market situations, since electricity prices typically experience significant fluctuations and frequent jumps, which are not captured well by a small volatility and low jump intensity.
@@
Figure 3: European call option in the Kou-type jump model. Total discretisation errors of the SLCNFI, CNFI and DIRKFI schemes for $N=\left\lceil{\frac{m}{2}}\right\rceil$ and set 4 (top), set 5 (middle), set 6 (bottom). Added: dashed reference line for convergence order 2.
@@
Figure 5: European call option in the Kou-type jump model. Temporal discretisation errors of the CNFI and DIRKFI schemes for $m=200$ and set 4 (top), set 5 (middle), set 6 (bottom). Added: dashed reference line for convergence order 2.
@@ 6.3 Numerical valuation of swing options
For the numerical valuation of swing options, we combine the numerical schemes proposed for the European call option case with the dynamic programming principle to find the optimal exercise at each action time. As before, we reverse the time to obtain an initial condition instead of a terminal condition. The action time $\tau_{n}$ in reversed time is $T-T_{N_{a}-n+1}$ ($n=1,2,\ldots,N_{a}$) and the sequence of PIDEs (2.1) then becomes
@@
The valuation procedure is outlined in Algorithm 1. At the first action time, $\tau_{1}=0$, the holder buys the maximal allowable amount. Given the option value at $\tau_{1}$, the PIDE is solved to obtain the option value up to the second action time, $\tau_{2}$. It is then considered whether it is optimal to exercise or not by maximising the option value at $\tau_{2}$. This process is repeated for each subsequent action time up to and including $\tau_{N_{a}}$. At each action time, the option value is determined by optimising over all feasible exercise amounts $b\in\{0,1,\ldots,L\}$, while ensuring that the cumulative purchased amount does not exceed the global constraint $M$. Finally, the PIDE is solved over the interval $(\tau_{N_{a}},\tau_{N_{a}+1}]$ to arrive at the desired swing option value function $v(\cdot,\cdot,0,T)$.
@@
6.4 Convergence behaviour: swing options
In this section, we study the total discretisation error of the three schemes from Section 4 in the numerical valuation of swing options. Here the number of swing action times is set equal to $N_{a}=20$, with local constraint given by $L=1$ and global constraint by $M=10$. Both Merton- and Kou-type jump models are considered, for the parameter sets 1 and 4 given by the Tables 1 and 2, respectively, with maturity time $T=1$. To assess the convergence rate of the schemes, we successively refine in this section the spatial and temporal grids, where at each refinement level the numbers of grid points in the $x$-, $y$- and $t$-directions are simultaneously doubled. Denote by $\widetilde{V}(N)$ the approximated swing option value at a given point $(x,y,z,T)$ if $m_{1}=m_{2}=N$, where $N$ is the number of time steps between each two consecutive action times. The numerical convergence rate is then defined by
$\text{rate}=\log_{2}\Bigg(\frac{|\widetilde{V}(N/2)-\widetilde{V}(N/4)|}{|\widetilde{V}(N)-\widetilde{V}(N/2)|}\Bigg).$
Table 3 reports for the two parameter sets 1 and 4 the approximated swing option values at three points $(x,y,0,T)$ using the CNFI scheme with $N=50,100,200,400$. For each set and point, the corresponding numerical convergence rate is added. Clearly, a favourable second-order convergence behaviour is observed. Numerical experiments in the case of the SLCNFI and DIRKFI schemes reveal the same positive result.
Table 3: Swing option values obtained with the CNFI scheme at several points $(x,y,z,T)$ with $z=0$ and $T=1$ for sets 1 and 4 where $N_{a}=20$, $L=1$, $M=10$. Numerical convergence rate added.
@@ 6.4 Convergence behaviour: swing options
$m_{1}=m_{2}=N$ $x=40,y=5$ rate x=60,y=100x=60,y=-100 rate $x=80,y=100$ rate
$50$ $681.8451$ $694.4583$ $710.0696$
$100$ $681.4723$ $694.1128$ $709.7014$
$200$ $681.3949$ $2.26$ $694.0408$ $2.26$ $709.6283$ $2.33$
$400$ $681.3740$ $1.89$ $694.0210$ $1.85$ $709.6078$ $1.85$
6.5 The Greeks and sensitivity analysis: swing options
The Greeks represent the sensitivities of the option value with respect to underlying variables and parameters, and are important for hedging purposes. They serve as an essential tool for quantifying the impact of changes in model inputs on the option value. In this section, we consider the Delta Greeks $\Delta_{1}=\frac{\partial v}{\partial x}$ and $\Delta_{2}=\frac{\partial v}{\partial y}$, which measure the sensitivity of the swing option value with respect to changes in the price components $x$ and $y$, respectively. Approximations to the Deltas are directly obtained at no additional computational cost, as they are naturally incorporated in the finite difference discretisation.
Set 1 Set 4
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Surfaces of the swing option value and its Delta Greeks if $z=0$ and $T=1$ for sets 1 and 4 where $N_{a}=20$, $L=1$, $M=10$. The left column displays the option value, $\Delta_{1}$ and $\Delta_{2}$ for set 1. The right column displays the option value, $\Delta_{1}$ and $\Delta_{2}$ for set 4.
Figure 6 displays the surfaces of the approximated swing option value together with its Delta Greeks if $z=0$ and $t=T$ for parameter sets 1 and 4 with the swing parameters $N_{a}=20$, $L=1$, $M=10$. A perusal of $\Delta_{1}$ indicates that three common regions in the $x$-domain can be distinguished. In the deep out-of-the-money region, $\Delta_{1}$ is positive but relatively close to zero, i.e., small changes in the price component $x$ have minor impact on the option value. Next, there is a region where $\Delta_{1}$ exhibits a sharp increase as $x$ increases, and a small change in $x$ has a significant impact on the option value. Finally, in the deep in-the-money region, $\Delta_{1}$ reaches a plateau, meaning that the option value increases linearly with $x$.
A perusal of $\Delta_{2}$ shows the same three regions. Notice that the sensitivity with respect to the price component $y$ – as captured by $\Delta_{2}$ – is significantly lower than that with respect to $x$, with about a factor $10^{-3}$. This reduced influence is primarily attributed to the high mean reversion speed $\beta$, which rapidly undermines the impact of $y$ over time. Nevertheless, it is anticipated that the influence of the price component $y$ will become more pronounced as the number of swing action times increases, since the option holder may benefit more frequently from price jumps before their effects are diminished by the strong mean-reverting behaviour.
6.6 Optimal exercise policy
In this subsection, we consider the optimal exercise policy for swing options and examine the influence of various model parameters, given in Table 4, on this policy. We select the Kou-type jump-diffusion model in view of its flexibility in independently modelling upward and downward jumps. The optimal exercise policy is obtained by applying Algorithm 2 after the approximation of the swing option values by the dynamic programming approach given in Algorithm 1.
Figure 7 presents a series of heat maps illustrating the cumulative amount of energy purchased up to forward time $t=0.5$ under a swing option contract with parameters $K=50$, $T=1$, $r=0.03$, $N_{a}=100$, $M=50$. The choice $t=0.5$ allows for a clear differentiation between three distinct cases in the optimal exercise policy. The first region (dark blue) represents no exercise, where no units have yet been bought up to time $t=0.5$. The second region (dark red) corresponds to full exercise, where the maximum allowable number of units $M$ has already been bought. The third region (from light blue to light red) represents partial exercise, where the contract holder has already bought a certain amount of electricity but can still buy additional units in the future in anticipation of more favourable market conditions.
Each heat map in Figure 7 corresponds to a specific parameter set listed in Table 4. Here set A is chosen as a reference set from which each of the other seven sets B, C,…, H differs by just one parameter value, allowing a comparative analysis of how individual parameters affect the optimal exercise policy. The $x$- and $y$-axes represent the two terms in the electricity price $x+y$ at initial time, i.e., today.
Table 4: Parameter sets for the Kou-type jump case. The time is measured in years.
parameters $\mu$ $\alpha$ $\beta$ $\sigma$ $\lambda$ $p$ $\eta_{1}$ $\eta_{2}$ $L$
Set A 80 8 126 11 52 0.5 0.01 0.01 1
Set B 80 8 126 11 52 0.9 0.01 0.01 1
Set C 80 8 126 11 52 0.1 0.01 0.01 1
Set D 80 8 126 11 10 0.5 0.01 0.01 1
Set E 80 8 10 11 52 0.5 0.01 0.01 1
Set F 80 1 126 11 52 0.5 0.01 0.01 1
Set G 80 8 126 11 52 0.5 0.01 0.01 2
Set H 80 8 126 11 52 0.5 0.01 0.01 5
Set A
Set B
Set C
Set D
Set E
Set F
Set G
Set H
Figure 7: Cumulative amounts of energy purchased up to forward time $t=0.5$ under a swing option contract with $K=50$, $T=1$, $r=0.03$, $N_{a}=100$, $M=50$ and the parameter sets given by Table 4.
Compared to set A, upward jumps are more likely for set B, since the value $p$ has increased to $p=0.9$. It is clear that the dark red region, corresponding to full exercise, is smaller compared to set A. This indicates that the holder of the swing option tends to delay exercising, anticipating higher future prices due to the increased probability of upward jumps.
Conversely, for set C downward jumps are more probable, as the value $p$ has decreased to $p=0.1$. Clearly, the dark red region is now larger compared to that of set A. This suggests that the holder tends to exercise more often, with less emphasis on waiting, as the opportunity to benefit from favourable future price movements is reduced.
Set D has less frequent jumps, since the value $\lambda$ has decreased to $\lambda=10$. Again, the dark red region is larger compared to that of set A, indicating that the holder tends to exercise more often as the probability of favourable future price movements is reduced.
Set E has a lower mean reversion speed for the $Y$ process, namely $\beta=10$. Here the dark red region is very small and the region in between dark red and dark blue, corresponding to partial exercise, is large. This suggests that the optimal exercise policy is to wait longer. This could be explained by two factors: (i) downward jumps have a more prolonged impact, discouraging immediate exercise due to sustained low prices, and (ii) favourable price conditions, once reached, tend to persist longer, making it more attractive to wait in anticipation of even better opportunities.
Set F has a lower mean reversion speed for the $X$ process, namely $\alpha=1$. As a distinctive feature, if the $x$-component of the price $x+y$ is high while $y<0$, then the optimal exercise policy tends to delay exercise, see the dark blue region at the bottom right in the subplot. Indeed, it may be argued that, when a downward jump occurs, the slow mean reversion ensures that the high price persists, thus reducing the need to exercise and allowing more time to wait for a more favourable, upward jump. Next, at the upper left in the subplot a swift transition is observed from dark blue (no exercise) to dark red (full exercise). We conjecture that this transition is still smooth.
Sets G and H have a larger number of units of energy that can be bought at each swing action time, with $L=2$ and $L=5$ respectively. Here the dark blue region expands. This is conceivable because the option holder can afford to wait longer when the price is low to moderate, in anticipation of more favourable future opportunities that allow buying a larger number of units.